OpenStudy (anonymous):

solve the polynomial inequality and graph the solution set on a number line. Express the solution set in interval notation (2x +1)(3x-4) > 0

OpenStudy (jiteshmeghwal9):

(2x+1)>0/3x-4 =(2x+1)>0 =2x>-1 =x>-1/2 & (3x-4)>0/2x+1 =3x-4>0 =3x>4 =x>4/3

OpenStudy (jiteshmeghwal9):